Aunt is 7 years younger than uncle and ages total 71. I was thinking A + B - 7 = 71. I added 7 to each side to make it A + B = 71 + 7. A + B = 78. divided 78 by two to make 39 and then subtracted the 7 from 39 to make 32. so one is 32 and one is 39 and it totals 71 but I am thinking that I am making more work than necessary?
this is a 4th grade problem and we are just starting.
thank you

Answer by ankor@dixie-net.com(22740)

Do it this way.
Let a = Aunt's age
then
(a+7) = Uncle's age
:
a + (a+7) = 71
2a = 71 - 7
2a = 64
a = 32 is Aunt's age
and
39 = uncles age
:
Check: 32 + 39 = 71